A family has four children find the probability that the family has equal number of sons and daughter.


1
Expert's answer
2022-06-09T10:58:47-0400

Let "X=" the number of boys:"X\sim Bin(n, p)."

Given "n=4, p=0.5, q=1-p=0.5"


"P(X=2)=\dbinom{4}{2}(0.5)^{2}(0.5)^{4-2}=\dfrac{3}{8}=0.375"